What is the mass of technical calcium carbide with a mass fraction of 12% required for the production of acetylene with a volume of 6.72 l (nu)?

1
Expert's answer
2022-02-17T05:57:05-0500

Solution:

calcium carbide = CaC2

acetylene = C2H2


At STP, one mole of any gas occupies a volume of 22.4 L.

Thus, 6.72 liters of C2H2 correspond to:

Moles of C2H2 = (6.72 L C2H2) × (1 mol C2H2 / 22.4 L C2H2) = 0.3 mol C2H2


Balanced chemical equation:

CaC2 + 2H2O → C2H2 + Ca(OH)2

According to stoichiometry:

1 mol of CaC2 produces 1 mol of C2H2

Thus, X mol of CaC2 produces 0.3 mol of C2H2

Moles of CaC2 = X = (0.3 mol C2H2) × (1 mol CaC2 / 1 mol C2H2) = 0.3 mol CaC2


The molar mass of CaC2 is 64.1 g/mol

Hence,

Mass of CaC2 = (0.3 mol CaC2) × (64.1 g CaC2 / 1 mol CaC2) = 19.23 g CaC2


Calculate the mass of technical CaC2:

Mass of technical CaC2 = (19.23 g CaC2) × (100% / 12%) = 160.25 g


Answer: The mass of technical calcium carbide (CaC2) is 160.25 grams